Mark put 3/4 of his earnings into his savings account.He spent the remaining money on treats for his dog.If he spent $36 on the treats,how much did Mark deposit?
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)

let deposit be x
3/4 x he saved
$36 he spent on treats
3/4 x +36 =x
3x+ 144=4x
x=144
he deposited $144

Answer by ikleyn(52908)

.
The correct answer is:

    Mark deposited 3*36 = 108 dollars.


Which is ABSOLUTELY OBVIOUS.


Be aware:  The answer by the other tutor is   W R O N G.